OpenStudy (anonymous):

Rewrite with positive exponents, and simplify, if possible. 9x/7^√y

OpenStudy (anonymous):

\[9x/7^\sqrt{y}\] is that the expression?

OpenStudy (anonymous):

\[9x / \sqrt[7]{y}]

OpenStudy (anonymous):

idk if i wrote that last one right but it is a little 7

OpenStudy (anonymous):

\[9x/\sqrt{7y}\] This?

OpenStudy (anonymous):

or: \[9x/\sqrt[7]{y}\]

OpenStudy (anonymous):

no everything is correct but it is a little 7 on top of the check

OpenStudy (kropot72):

\[\frac{9x}{\sqrt[7]{y}}\]

OpenStudy (anonymous):

yes thats it

OpenStudy (anonymous):

I'm not exactly sure what this question is asking, except perhaps that they want you to change \[\sqrt[7]{y}\] into exponent form. When you have the square root, you can rewrite it in exponent form like this: \[\sqrt[2]{y} = y ^{1/2}\] And when you take the inverse of an exponent, you change the sign on the exponent: \[1 / y ^{1/2} = y ^{-1/2}\] So you can convert \[\sqrt[7]{y}\] into an exponent. Perhaps that's all?

OpenStudy (anonymous):

I don't think thats right. the answer choices i have are 9xy^1/7 or 9x/y^7 or 9xy^7 or 9x/y^1/7

OpenStudy (anonymous):

Ok, so just do the first thing -- convert \[\sqrt[7]{y}\] into an exponent like I did the example of \[\sqrt[2]{y}=y ^{1/2}\]